# On Bayes factor functions

### v2

@@ Abstract
Bayes Factor Functions (BFFs) represent Bayes factors as functions of the hyperparameters indexing the prior distributions used to define alternative hypotheses (Johnson, Pramanik and Shudde, 2023; Datta et al., 2025). While earlier work has primarily focused on normal-moment and gamma priors, we build on recent results by Wagenmakers and Grasman (2025) to explore broader classes of priors for constructing BFFs. For tests based on common test statistics, our findings show that inverse-moment and inverse-gamma priors can yield higher expected posterior probabilities in favor of both true null and true alternative hypotheses than other default priors. Using test statistics from replicated experiments, we also demonstrate how method-of-moments estimators can inform prior parameter selection within these families, thereby enhancing inference in replicated studies and meta-analyses.
62F15,
62F03,
62F05,
Bayes factor based on test statistic,
Meta-analysis,
Non-local prior density,
@@ 1 Introduction
Bayes factors provide a compelling alternative to $p$-values for summarizing the outcomes of hypothesis tests. In contrast to classical significance testing, Bayesian methods explicitly quantify the strength of evidence in favor of both the null and alternative hypotheses, offering a coherent interpretation based on the relative plausibility of the observed data under each model. However, the validity and interpretability of Bayesian test results are inherently tied to the choice of prior distributions—either a specific prior in the case of Bayes factors, or a class of priors when using Bayes factor functions (BFFs).
The need to specify prior densities on parameters that define alternative hypotheses (i.e., alternative prior densities), as well as prior probabilities for the hypotheses themselves, has limited the widespread adoption of Bayesian tests. When testing only two hypotheses, the latter issue is often sidestepped by assigning equal prior probability to each hypothesis or by simply reporting a Bayes factor. Nonetheless, defining alternative prior densities remains a significant challenge.
To address the first issue, we advocate the report of Bayes factor functions (BFFs) (Johnson, Pramanik and Shudde, 2023). Instead of reflecting the value of a single Bayes factor, BFFs provide a summary of a continuum of Bayes factors obtained from various classes of alternative prior densities.
As proposed in (Johnson, Pramanik and Shudde, 2023), BFFs were defined from the sampling distributions of classical test statistics with alternative hypotheses defined using non-local moment prior and gamma prior densities. In this article, we extend the application of BFFs in the following ways:
  • In Section 2, we formally define BFFs and illustrate their construction. Inverse-moment densities are introduced for tests of normal means, and BFFs are constructed using these densities. The BFF constructed using inverse-moment priors is informally compared to a BFF constructed using Cauchy distributions.
  • Section 3 uses an equality exposed in Wagenmakers and Grasman (2025) to compare the expected posterior probabilities of Bayes factors and BFFs based on various classes of prior densities, including inverse-moment, g-prior, Cauchy, normal, gamma, and inverse-gamma prior densities.
  • Section 4 discusses the selection of the shape and scale parameters for inverse-moment and inverse gamma prior densities and describes an interpretable procedure for indexing BFFs according to standardized effect sizes.
  • Examples that illustrate the construction of BFFs are provided in Section 5.
  • Section 6 presents an analysis of replicated experiments using Bayes factor functions (BFFs) and demonstrates how non-local alternative priors naturally induce a hierarchical structure on effect sizes across studies. This structure arises by scaling the noncentrality parameters of the test statistics in each experiment according to their respective sample sizes, thereby allowing the assumption that these parameters are independently drawn from priors centered on a common underlying effect size.
@@
The BFFs described in this article are defined from the sampling distributions of test statistics rather than full data likelihoods. We adopt this approach for three reasons. First, it simplifies the specification of Bayesian hypothesis tests because the sampling distributions of test statistics under null hypotheses are known, and the sampling distributions of test statistics under alternative hypotheses often depend only on scalar noncentrality parameters. This fact simplifies the specification of priors necessary for a Bayesian hypothesis test. Second, many–if not most–hypothesis tests reported in the social sciences, biological sciences, and medical sciences are based on classical test statistics. Exposing methods to easily construct BFFs using results from these tests promises to significantly expand the application of Bayesian testing methods in these fields. Finally, constructing BFFs from test statistics makes it possible to easily aggregate information from published research studies in systematic reviews and meta-analyses.
@@
BFFs can also be defined using full data likelihood functions and prior densities specified on parameters of arbitrary dimension. In such cases, Bayes factor surfaces can also be used to reflect the variation of Bayes factors as a function of prior hyperparameters (Franck and Gramacy, 2020). However, interpreting these surfaces in common testing scenarios, especially in null hypothesis significance tests, can be challenging. It is also more difficult to aggregate information from Bayes factor curves for replicated experiments. A critical feature of the methodology described below is the expression of Bayes factors as interpretable functions of the discrepancy between null and alternative hypotheses.
@@ 2 Construction of BFFs
To illustrate the construction of a BFF based on a test statistic $x$, we assume the following parametric Bayesian framework inspired by null hypothesis significance testing. Specifically, we define the hypotheses as follows:
$H_{0}$ $:$ $x\sim f(\cdot\mid\lambda_{0}),\qquad\lambda_{0}\mbox{ known},$ (1)
$H_{1}$ $:$ $x\sim f(\cdot\mid\lambda),\qquad\lambda\sim g(\cdot\mid\psi).$ (2)
@@
$BF_{10}(x{\,|\,}\psi)=\frac{\int_{\Lambda}f(x\mid\lambda)g(\lambda\mid\psi)d\lambda}{f(x\mid\lambda_{0})},$ (4)
@@
To illustrate these ideas, let us construct a BFF for a two-sided $z$-test where the data $x_{1},\dots,x_{n}$ are independently distributed as $N(\mu,\sigma^{2})$ with $\sigma^{2}$ known. The test statistic $z=\sqrt{n}x/\sigma$ follows a $N(\sqrt{n}\omega,1)$ distribution, where $\omega:=\mu/\sigma$ represents the standardized effect. The null hypothesis is $\mu=0$, or equivalently $\omega=0$. For the alternative hypothesis, instead of specifying a prior on $\mu$, we specify a prior on the non-centrality parameter $\lambda:=\sqrt{n}\omega$.
For illustration, we assume an inverse-moment prior density for $\lambda$. That is, we assume that $\lambda$ is drawn from an inverse-moment prior density ${i}(\lambda{\,|\,}\tau,\nu)$ of the form (Johnson and Rossell, 2010)
$i(\lambda{\,|\,}\tau,\nu)=\frac{\tau^{\frac{\nu}{2}}}{\Gamma\left(\frac{\nu}{2}\right)}(\lambda^{2})^{-\frac{\nu+1}{2}}\exp\left(-\frac{\tau}{\lambda^{2}}\right),\ -\infty<\lambda<\infty,\quad\tau,\nu>0.$ (5)
The modes of this density occur at
@@
For one-sided tests, the densities $i^{+}(\cdot{\,|\,},\tau,\nu)$ and $i^{-}(\cdot{\,|\,},\tau,\nu)$ are defined as $2i(\cdot{\,|\,},\tau,\nu)$, but constrained to either the positive or negative real lines, respectively. Fig. 1 shows several examples of inverse-moment densities. We denote the corresponding distribution function by $\mathcal{I}(\lambda{\,|\,}\tau,\nu)$.
@@
Given fixed values of $(\tau,\nu)$, a hypothesis test can then be specified as
@@
Under the null hypothesis, the marginal density of $z$, $m_{0}(z)$, is a standard normal density, while under the alternative hypothesis, the marginal density of $z$ is
$m_{1}(z\mid\tau,\nu)=\int_{-\infty}^{\infty}\frac{1}{\sqrt{2\pi}}\exp\left[-\frac{(z-\lambda)^{2}}{2}\right]\frac{\tau^{\frac{\nu}{2}}}{\Gamma\left(\frac{\nu}{2}\right)}(\lambda^{2})^{-\frac{\nu+1}{2}}\exp\left(-\frac{\tau}{\lambda^{2}}\right)\,d\lambda.$ (7)
Although this one-dimensional integral cannot be evaluated analytically, it can be evaluated numerically in many statistical packages, including R.
We denote the Bayes factor in favor of $H_{1}$ by $BF_{10}(z{\,|\,}\tau,\nu)=m_{1}(z\mid\tau,\nu)/m_{0}(z)$. The logarithm of the Bayes factor in favor of the alternative, $\log BF_{10}(z{\,|\,}\tau,\nu)$, is called the weight of evidence, WOE.
@@
Thus, the Bayes factor can be expressed as a function of the standardized effect $\omega_{1}$, and a BFF can be defined as the mapping of standardized effects to Bayes factors obtained as $\omega_{1}$ is varied over its domain.
Criteria for selecting $\nu$ for this particular class of priors are discussed below.
@@
Figure 2: Prior densities and BFFs. (a) Inverse moment prior density for $\nu=9$, $\omega=0.5$, $n=100$. This density has its modes at $\lambda=\pm\sqrt{n}\omega=\pm 5$. (b) BFF generated from inverse moment prior density, on logarithmic scale. This plot displays strong evidence (i.e., $\log[BF_{10}(z)]<-3$) against alternative prior densities corresponding to $\omega_{1}>0.35$, and very strong evidence (<5<-5) for $\omega_{1}>0.46$. (c) Cauchy prior on noncentrality parameter with scale $r=5$. (d) BFF based on Cauchy prior as scale parameter $r$ is increased.
Several features of Fig. 2(b) deserve mention. With $z=1$, the BFF is a decreasing function of $\omega_{1}$. The BFF demonstrates positive support for the null hypothesis for inverse-moment priors centered on $\omega_{1}>0.12$ ($\mbox{WOE}<-1$), strong support for $\omega_{1}>0.32$ ($\mbox{WOE}<-3$), and very strong support for $\omega_{1}>0.50$ ($\mbox{WOE}<-5$).
This figure also shows that the BFF converges to 0 as $\omega_{1}$ goes to 0. This property holds for all values of the $z$ statistic because the inverse-moment prior densities centered on small $\omega$ values concentrate their mass near the null parameter value, resulting in approximately the same marginal probability being assigned to the data under both hypotheses.
The construction of BFFs for $t$, $\chi^{2}$, and $F$ tests proceeds similarly. The inverse-moment prior density given in (5) provides a flexible default alternative prior density for one- and two-sided, one- and two-sample $z$ and $t$ tests. For $\chi_{k}^{2}$ and $F_{k,m}$ tests, inverse-gamma prior densities, $IG(\frac{\nu}{2},\tau)$, may be placed on the non-centrality parameter. Notably, if $W\sim\mathcal{I}(\nu,\tau)$, then $W^{2}\sim IG(\frac{\nu}{2},\tau)$, establishing a natural connection between these prior choices.
Under these prior choices, the Bayes factor based on a two-sided $z$ statistic equals the Bayes factor based on the $\chi^{2}_{1}$ statistic $z^{2}$, and the Bayes factor based on a two-sided $t_{k}$ statistic equals the Bayes factor based on the $F_{1,k}$ statistic $t_{k}^{2}$, provided the same values of $\nu$ and $\tau$ are used to define the tests.
While inverse-moment and inverse-gamma priors serve as convenient defaults, BFFs can also be defined using other prior distributions for the non-centrality parameters. For example, Fig 2(c) and (d) illustrate the definition of a BFF using a Cauchy prior for the $z$-test described above (e.g., Rouder et al., 2009, 2012). The scale parameter $r$ of the Cauchy prior in Fig. 2(c) is five, and its interquartile range 10, equaling the distance between the modes of the inverse-moment prior in Fig. 2(a). The tail behaviors of the two prior densities are similar. The BFF based on the Cauchy prior depicted in Fig. 2(d) provides “positive” evidence in favor of the null hypothesis for priors with $r>3$, and “strong evidence” for $r>26$.
@@ 3 Comparisons of alternative classes of prior distributions
The asymptotic behavior of Bayes factors under true point null hypotheses is largely determined by the properties of the alternative prior density at the null value of the parameter being tested (Johnson and Rossell, 2010). In the present framework, hypothesis tests assess whether the noncentrality parameter $\lambda$ of a test statistic equals zero. Consequently, the behavior of the prior distribution for $\lambda$ near zero directly affects the properties of the Bayes factor.
@@
sponding to a continuous local alternative—and under certain regularity conditions, the Bayes factor in favor of the alternative hypothesis, when the null is true, converges at the rate $BF_{10}=O_{p}(n^{-1/2})$ (Johnson and Rossell, 2010). For specific test statistics, such as the $z$, $t$, $\chi^{2}$, and $F$ statistics, Datta et al. (2025) derived closed-form expressions for Bayes factors based on normal moment priors. Their results show that if the normal moment prior density behaves as $O((\lambda^{2})^{r})$ as $\lambda\to 0$, then the Bayes factor in favor of a true null hypothesis satisfies $BF_{10}=O_{p}(n^{-r-1/2})$ for $z$ and $t$ tests, and $BF_{10}=O_{p}(n^{-r-k/2})$ for $\chi^{2}_{k}$ and $F_{k,m}$ tests.
In contrast, if inverse-moment priors are assumed for the non-centrality parameters of $z$ and $t$ tests, and inverse gamma (IG) priors are imposed on noncentrality parameters for $\chi^{2}$ tests, the convergence of resulting Bayes factors in favor of true null hypotheses is of exponential order in $\sqrt{n}$. Specifically, the following theorems apply.
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<1.$ (8)
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<1.$ (9)
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<1.$ (10)
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<1.$ (11)
@@
For true alternative hypotheses, $BF_{01}$ is typically $O_{p}(\exp(-cn))$ for some $c>0$ for all alternative priors that are continuous and positive at the (true) non-null parameter (Bahadur and Bickel, 2009).
In the case of finite samples, of course, asymptotic convergence rates can be misleading, and practical performance may differ substantially from asymptotic expectations.
To evaluate the finite sample performance of Bayes factors, we draw on a recent result of Wagenmakers and Grasman, who exposed the following equality. Here, $f$ and $g$ are prior predictive distributions for two models, ${\cal{M}}_{f}$ and ${\cal{M}}_{g}$, each assigned equal prior probability. Then from Wagenmakers and Grasman (2025),
$\int f(x)\frac{f(x)}{f(x)+g(x)}dx=\int g(x)\frac{g(x)}{f(x)+g(x)}dx.$ (12)
With prior equipoise, this equality states that the expected posterior probability (EPP) of ${\cal{M}}_{f}$ when ${\cal{M}}_{f}$ is true equals that of ${\cal{M}}_{g}$ when ${\cal{M}}_{g}$ is true. We denote the value of this discrepancy measure by ${\rm D_{EP}}(f\parallel g)$.
We can use this measure to compare BFFs defined by different alternative prior distributions. Specifically, letting ${\cal M}_{0}$ denote the null model $H_{0}$, and ${\cal M}_{1a}$ and ${\cal M}_{1b}$ denote two alternative models indexed by a common parameter $\xi$, we compare ${\rm D_{EP}}(0\parallel 1a(\xi))$ to ${\rm D_{EP}}(0\parallel 1b(\xi))$. Larger values of $\rm D_{EP}$} imply higher expected posterior probability support of the true model, whether it is the null or alternative.
When comparisons are restricted to non-local priors, the prior modes provide a natural indexing parameter. However, local alternative priors used in standard Bayesian null hypothesis significance tests are often centered on the null value, making the mode unsuitable for indexing. Additionally, certain priors–such as the Cauchy–do not possess finite moments and so cannot be indexed by means or standard deviations. To accommodate these cases, alternative priors can be indexed by their interquartile range (IQR) in two-sided tests and by their medians in one-sided tests. This parameterization provides an interpretable basis for comparing the performance of BFFs across a range of prior distributions.
To compare choices of alternative prior densities using this methodology, we begin by considering BFFs based on $t$-tests. The null hypothesis is that the non-centrality parameter of the test statistic satisfies $\lambda=0$. The following alternative choices of prior densities on the non-centrality parameter are compared:
@@
    An inverse-moment prior $i(\lambda{\,|\,}\tau,\nu)$ with shape parameter $\nu=1$.
@@
    A Cauchy prior $C(\lambda{\,|\,}0,r)$, where $r$ is the scale parameter (e.g., Jeffreys, 1961; Zellner and Siow, 1984, 1980) .
@@
    A normal prior, $N(\lambda{\,|\,}0,g)$.

### v3

@@ Abstract
Bayes Factor Functions (BFFs) represent Bayes factors as functions of prior hyperparameters (Johnson et al.,, 2023; Datta et al.,, 2025). While earlier work has primarily focused on normal-moment and gamma priors, we build on recent results by Wagenmakers and Grasman, (2025) to explore broader classes of priors for constructing BFFs. For tests based on common test statistics, our findings show that inverse-moment and inverse-gamma priors can yield higher expected posterior probabilities in favor of both true null and true alternative hypotheses than other default priors. Using test statistics from replicated experiments, we also demonstrate how method-of-moments estimators can inform prior parameter selection within these families, thereby enhancing inference in replicated studies and meta-analyses.
@@ 1 Introduction
Bayes factors provide a compelling alternative to $p$-values for summarizing hypothesis test outcomes. Unlike classical significance testing, they quantify evidence for both null and alternative hypotheses, yielding an interpretable measure of how plausible the observed data are under competing models. Introduced by Jeffreys in 1935, Bayes factors are defined as ratios of marginal likelihoods that directly measure evidential weight in hypothesis testing (Jeffreys,, 1935, 1939) [^1] Haldane (1932) had already introduced a mixture prior formulation that anticipated similar ideas (Haldane,, 1932; Etz and Wagenmakers,, 2017). Building on Jeffreys’ foundation, Berger, (1985) established their role in Bayesian model choice and clarified their link to posterior model probabilities. Later, Kass and Raftery, (1995) described computational strategies and refined interpretive guidelines, further encouraging their application.
In the last two decades, many have argued that Bayes factors should replace or supplement $p$-values as the primary measure of evidence against a null hypothesis (e.g., Bayarri and Berger,, 2004; Johnson,, 2013; Held,, 2018; Rougier,, 2019; Benjamin et al.,, 2018; Benjamin and Berger,, 2019). Nonetheless, their adoption has been limited by the need to specify both prior densities on model parameters and prior probabilities for the hypotheses themselves. When only two hypotheses are compared, the latter issue is often avoided by assigning equal prior weights or simply reporting the Bayes factor. However, defining suitable alternative priors remains a central challenge, and numerous proposals have been developed.
Jeffreys (1939) was the first to suggest a Cauchy prior for testing whether a normal mean equals zero, laying the groundwork for the Jeffreys–Zellner–Siow (JZS) prior (Zellner and Siow,, 1980, 1984). The JZS prior, based on a scaled Cauchy distribution, was later operationalized by Rouder et al., (2009) and incorporated in the BayesFactor package by Morey and Rouder, (2023). Other important contributions include intrinsic priors (Berger and Pericchi,, 1996, 1996; Moreno et al., 1998b, ; Moreno et al., 1998a, ; Casella and Moreno,, 2006; Berger et al.,, 2001), which require specification of a minimal training sample size, and fractional Bayes factors (O’Hagan,, 1995, 1997), which depend on selecting a training fraction of the data. Conjugate $g$-priors (e.g., Zellner and Siow,, 1984; Zellner,, 1986; Liang et al.,, 2008) offer another approach, later extended to generalized linear models by Bové and Held, (2011). Bayarri et al., 2012a and Bayarri et al., 2012b provide criteria that objective priors should satisfy and show that many previously proposed priors, with modification, do. Nonlocal priors represent yet another class. They vanish at parameter values consistent with the null hypothesis and yield faster accumulation of evidence in favor of both true null and true alternative hypotheses (Johnson and Rossell,, 2010), but they also require specification of scale parameters. A comprehensive overview of these and other objective priors is provided by Consonni et al., (2018).
A persistent difficulty across these approaches is the selection of hyperparameters that determine prior scale. To address this problem, we advocate reporting Bayes factor functions (BFFs) (Johnson et al.,, 2023). Rather than presenting a single Bayes factor, BFFs summarize the continuum of Bayes factors obtained from a class of alternative priors.
@@
In this article, we focus on BFFs defined using the sampling distributions of test statistics rather than full data likelihoods (Johnson,, 2005). This choice simplifies matters in several ways. First, the null distributions of classical test statistics are known, and their alternatives often depend only on a scalar noncentrality parameter, which simplifies prior specification. Second, most hypothesis tests reported in the social, biological, and medical sciences are based on classical test statistics, making this approach widely applicable (e.g. Skaik,, 2015; Liu and Wang,, 2021). Third, constructing BFFs from test statistics facilitates the aggregation of evidence across studies, enabling systematic reviews and meta-analyses.
@@
Of course, BFFs can also be defined using full likelihoods and multidimensional parameter priors. In such cases, Bayes factor surfaces can be used to display variation as a function of prior hyperparameters (Franck and Gramacy,, 2020). However, these are often more difficult to interpret in routine testing settings and less convenient for aggregating replicated results. A critical feature of our approach is that Bayes factors are expressed as interpretable functions of the discrepancy between null and alternative hypotheses.
The remainder of this paper proceeds as follows:
  • Section 2 formally defines BFFs, illustrating their construction with inverse-moment priors for tests of normal means, and compares them informally to BFFs based on Cauchy priors (Jeffreys,, 1961; Zellner and Siow,, 1980).
  • Section 3 applies an equality from Wagenmakers and Grasman, (2025) to compare expected posterior probabilities across different prior classes, including inverse-moment, $g$-prior, Cauchy, normal, gamma, and inverse-gamma densities.
  • Section 4 addresses the choice of shape and scale parameters for inverse-moment and inverse-gamma priors and proposes a procedure for indexing BFFs by standardized effect sizes.
  • Section 5 presents worked examples demonstrating the construction of BFFs.
  • Section 6 analyzes replicated experiments, showing how nonlocal priors induce a natural hierarchical structure on effect sizes across studies by scaling noncentrality parameters according to sample size, thereby allowing them to be modeled as draws from a common prior on effect size.
@@ 2 Construction of BFFs
To illustrate the construction of a BFF based on a test statistic $x$, we assume the following parametric Bayesian framework inspired by null hypothesis significance testing (see Ly et al., (2007) and comments in Petrone, (2007) for connections to interval null hypothesis tests). Specifically, we define the hypotheses as follows:
$H_{0}$ $:$ $x\sim f(\cdot\mid\lambda_{0}),\qquad\lambda_{0}\ \mbox{is known},$
$H_{1}$ $:$ $x\sim f(\cdot\mid\lambda),\qquad\lambda\sim g(\cdot\mid\psi).$
@@
$BF_{10}(x{\,|\,}\psi)=\frac{\int_{\Lambda}f(x\mid\lambda)g(\lambda\mid\psi)d\lambda}{f(x\mid\lambda_{0})},$
@@
To illustrate the construction of a BFF, consider a two-sided $z$-test with data $x_{1},\dots,x_{n}$ independently distributed as $N(\mu,\sigma^{2})$, where $\sigma^{2}$ is known. The standardized test statistic is $z=\sqrt{n}x/\sigma$, which follows a $N(\sqrt{n}\omega,1)$ distribution, with $\omega:=\mu/\sigma$ denoting the standardized effect. The null hypothesis specifies $\mu=0$, or equivalently $\omega=0$. For the alternative, rather than placing a prior directly on $\mu$, we specify a prior on the noncentrality parameter $\lambda:=\sqrt{n}\omega$.
As a default prior for $\lambda$, we adopt the inverse-moment distribution (Johnson and Rossell,, 2010), defined by
$i(\lambda\mid\tau,\nu)=\frac{\tau^{\nu/2}}{\Gamma(\nu/2)}(\lambda^{2})^{-(\nu+1)/2}\exp\left(-\frac{\tau}{\lambda^{2}}\right),\quad-\infty<\lambda<\infty,\ \tau,\nu>0.$ (1)
This density has modes at
@@
For one-sided alternatives, the densities $i^{+}(\cdot\mid\tau,\nu)$ and $i^{-}(\cdot\mid\tau,\nu)$ are defined as
$2i(\cdot\mid\tau,\nu)$, restricted to the positive or negative half-line, respectively. Several examples of these priors are shown in Fig. 1. We denote the corresponding distribution by $\mathcal{I}(\lambda\mid\tau,\nu)$.
@@
With $(\tau,\nu)$ fixed, the hypotheses can be written as
@@
Under $H_{0}$, the marginal density of $z$ is standard normal. Under $H_{1}$, the marginal density is
$m_{1}(z\mid\tau,\nu)=\int_{-\infty}^{\infty}\frac{1}{\sqrt{2\pi}}\exp\left[-\frac{(z-\lambda)^{2}}{2}\right]\frac{\tau^{\frac{\nu}{2}}}{\Gamma\left(\frac{\nu}{2}\right)}(\lambda^{2})^{-\frac{\nu+1}{2}}\exp\left(-\frac{\tau}{\lambda^{2}}\right)\,d\lambda.$
Although this integral lacks a closed form, it can be readily computed numerically in software such as R, and its logarithm, $\log BF_{10}(z\mid\tau,\nu)$, is called the weight of evidence (WOE) (Good,, 1985; Kass and Raftery,, 1995).
@@
Thus, the Bayes factor can be expressed as a function of the standardized effect $\omega_{1}$, and a BFF can be defined as the mapping of standardized effects to Bayes factors obtained as $\omega_{1}$ is varied over its domain. Criteria for selecting $\nu$ for this particular class of priors are discussed in Section 4.
@@
Figure 2: Prior densities and BFFs. (a) Inverse moment prior density for $\nu=9$, $\omega_{1}=0.5$, $n=100$. This density has its modes at $\lambda=\pm\sqrt{n}\omega_{1}=\pm 5$. (b) BFF generated from inverse moment prior density, on logarithmic scale. This plot displays strong evidence (i.e., $\log[BF_{10}(z)]<-3$) against alternative prior densities corresponding to $\omega_{1}>0.32$, and very strong evidence (<5<-5) for $\omega_{1}>0.50$. (c) Cauchy prior on noncentrality parameter with scale $r=5$. (d) BFF based on Cauchy prior as scale parameter $r$ is increased.
Several features of Fig. 2(b) deserve mention. With $z=1$, the BFF is a decreasing function of $\omega_{1}$. The BFF demonstrates positive support for the null hypothesis for inverse-moment priors centered on $\omega_{1}>0.12$ ($\mbox{WOE}<-1$), strong support for $\omega_{1}>0.32$ ($\mbox{WOE}<-3$), and very strong support for $\omega_{1}>0.50$ ($\mbox{WOE}<-5$). This figure also shows that the log(BFF) converges to 0 as $\omega_{1}$ goes to 0. This property holds for all values of the $z$ statistic because the inverse-moment prior densities centered on small $\omega$ values concentrate their mass near the null parameter value, resulting in approximately the same marginal probability being assigned to the data under both hypotheses.
The same construction extends to $t$, $\chi^{2}$, and $F$ tests. For $\chi^{2}_{k}$ and $F_{k,m}$ statistics, inverse-gamma priors $IG(\nu/2,\tau)$ on the noncentrality parameter provide a natural analogue, as $W\sim\mathcal{I}(\nu,\tau)$ implies $W^{2}\sim IG(\nu/2,\tau)$. Under these priors, the Bayes factor for a two-sided $z$ statistic equals that for the $\chi^{2}_{1}$ statistic $z^{2}$, and likewise for a two-sided $t_{k}$ and its corresponding $F_{1,k}$ statistic (provided the same values of $\nu$ and $\tau$ are used to define the tests).
While inverse-moment and inverse-gamma priors serve as convenient defaults, BFFs can also be defined using other prior distributions for the non-centrality parameters. For example, Fig 2(c) and (d) illustrate the definition of a BFF using a Cauchy prior for the $z$-test described above (e.g., Rouder et al.,, 2009, 2012). The scale parameter $r$ of the Cauchy prior in Fig. 2(c) is five, and its interquartile range 10, equaling the distance between the modes of the inverse-moment prior in Fig. 2(a). The tail behaviors of the two prior densities are similar. The BFF based on the Cauchy prior depicted in Fig. 2(d) provides “positive” evidence in favor of the null hypothesis for priors with $r>3$, and “strong evidence” for $r>26$.
@@ 3 Comparisons of alternative classes of prior distributions
The asymptotic behavior of Bayes factors under true point null hypotheses is largely determined by the properties of the alternative prior density at the null value of the parameter being tested (Johnson and Rossell,, 2010). In the present framework, hypothesis tests assess whether the noncentrality parameter $\lambda$ of a test statistic equals zero. Consequently, the behavior of the prior distribution for $\lambda$ near zero directly affects the properties of the Bayes factor.
@@
sponding to a continuous local alternative—and under certain regularity conditions, the Bayes factor in favor of the alternative hypothesis, when the null is true, converges at the rate $BF_{10}=O_{p}(n^{-1/2})$ (Johnson and Rossell,, 2010). For specific test statistics, such as the $z$, $t$, $\chi^{2}$, and $F$ statistics, Datta et al., (2025) derived closed-form expressions for Bayes factors based on normal moment priors of the form
$p(\lambda)=\frac{(\lambda^{2})^{r}}{(2\tau^{2})^{r+{\frac{1}{2}}}\Gamma(r+{\frac{1}{2}})}\exp\left(-\frac{\lambda^{2}}{2\tau^{2}}\right).$ (3)
imposed on noncentrality parameters $\lambda$ of $z$ and $t$ tests, and $Gamma(\lambda{\,|\,}k/2+r,1/(2\tau^{2})$ priors for $\chi^{2}_{k}$ and $F_{k,m}$ tests. Their results show that Bayes factors in favor of a true null hypothesis satisfies $BF_{10}=O_{p}(n^{-r-1/2})$ for $z$ and $t$ tests, and $BF_{10}=O_{p}(n^{-r-k/2})$ for $\chi^{2}_{k}$ and $F_{k,m}$ tests.
In contrast, if inverse-moment priors are assumed for the non-centrality parameters of $z$ and $t$ tests, and inverse gamma (IG) priors are imposed on noncentrality parameters for $\chi^{2}$ and $F$ tests, the convergence of resulting Bayes factors in favor of true null hypotheses is of exponential order in $\sqrt{n}$. Specifically, the following theorems apply.
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<{\frac{1}{2}}.$
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<{\frac{1}{2}}.$
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<{\frac{1}{2}}.$
@@
$BF_{10}(x)=O_{p}(\exp\left[-c\sqrt{n}\right])\quad\mbox{for some}\quad 0<c<{\frac{1}{2}}.$
@@
For true alternative hypotheses, $BF_{01}$ is typically $O_{p}(\exp(-cn))$ for some $c>0$ for all alternative priors that are continuous and positive at the (true) non-null parameter (Bahadur and Bickel,, 2009). In the case of finite samples, of course, asymptotic convergence rates can be misleading, and practical performance may differ substantially from asymptotic expectations.
To evaluate the finite sample performance of Bayes factors, we draw on a recent result of Wagenmakers and Grasman, who exposed the following equality. Here, $f$ and $g$ are prior predictive distributions for two models, ${\cal{M}}_{f}$ and ${\cal{M}}_{g}$, each assigned equal prior probability. Then Wagenmakers and Grasman, (2025) define the discrepancy measure
${\rm D_{EP}}(f\parallel g)=\int f(x)\frac{f(x)}{f(x)+g(x)}dx=\int g(x)\frac{g(x)}{f(x)+g(x)}dx,$
with ${\rm D_{EP}}(f\parallel g)={\rm D_{EP}}(g\parallel f)$. When equal prior probability is assigned to the two models, this equality states that the expected posterior probability of ${\cal{M}}_{f}$ when ${\cal{M}}_{f}$ is true equals the expected posterior probability of ${\cal{M}}_{g}$ when ${\cal{M}}_{g}$ is true. We write $\rm D_{EP}$} as shorthand for ${\rm D_{EP}}(f\parallel g)$ when the compared models are clear from context.
We can use this measure to compare BFFs defined by different alternative prior distributions. Specifically, letting ${\cal M}_{0}$ denote the null model $H_{0}$, and ${\cal M}_{1a}$ and ${\cal M}_{1b}$ denote two alternative models, we compare ${\rm D_{EP}}(0\parallel 1a)$ to ${\rm D_{EP}}(0\parallel 1b)$. If ${\rm D_{EP}}(0\parallel 1a)>{\rm D_{EP}}(0\parallel 1b)$, then the expected posterior probability of the true model when the null is compared to ${\cal M}_{1a}$ is higher than the expected posterior probability of the true model when the null is compared to ${\cal M}_{1b}$, assuming equal prior probabilities are assigned to each model in both comparisons.
To compare the efficiency of non-local priors in accumulating evidence against a null hypothesis, prior modes provide a natural indexing parameter. However, local alternative priors used in standard Bayesian null hypothesis significance tests are often centered on the null value, making the mode unsuitable for indexing. Additionally, certain priors–such as the Cauchy–do not possess finite moments and so cannot be indexed by means or standard deviations. To accommodate these cases, alternative priors can be indexed by their interquartile range (IQR) in two-sided tests and by their medians in one-sided tests. This parameterization provides an interpretable basis for comparing the performance of BFFs across a range of prior distributions.
To compare choices of alternative prior densities using this methodology, we begin by considering BFFs based on $t$-tests. The null hypothesis is that the non-centrality parameter of the test statistic satisfies $\lambda=0$. The following alternative classes of prior densities on the non-centrality parameter are compared:
@@
    Inverse-moment priors $i(\lambda{\,|\,}\tau,\nu)$ with shape parameter $\nu=1$.
@@
    Cauchy priors $C(\lambda{\,|\,}0,r)$, where $r$ is the scale parameter (e.g., Jeffreys,, 1961; Rouder et al.,, 2009).
@@
    Normal priors, $N(\lambda{\,|\,}0,g)$ (e.g., Zellner and Siow,, 1980, 1984).